# CoverUp: Coverage-Guided LLM-Based Test Generation Juan Altmayer Pizzorno

### v2

@@ 3.2. Initial Prompting
Figure 3 shows an example of an initial prompt, with circled numbers
identifying sections.
@@
a statement assigning the LLM the persona (White et al., 2023)
of an “expert Python test-driven developer”, intended to help guide it towards
high quality tests;
a sentence pointing out the code excerpt (segment), identifying what
file it comes from, stating what lines or branches do not execute when tested.
The portion specifying the lines and branches missing coverage is compressed
using line ranges, simplifying the prompt and reducing token usage.
a request for a pytest test functions, and an
encouragement for the LLM to use the provided tool function;
a series of other requests, such as
“include assertions” and “avoid state pollution”, to steer the result towards usable tests;
a request that the response only include the new Python tests
to facilitate extracting it from the response and to reduce token usage; and
@@
We find that tests generated by GPT-4 often include top-level code calling
into pytest.main or into parts of the test itself.
While such top-level code can make sense in a standalone test file, Python
executes it as part of the loading process and doing so may greatly
disrupt pytest’s operation.
In fact, the calls in some of these early tests caused pytest to restart
its test discovery, slowing it down to the point that it became unusable.
For that reason, part 4 of the prompt directs the LLM to
not include such calls.
Prompt part 4 also instructs the LLM to generate tests that
“clean up […], so as to avoid state pollution”.
It suggests the use of monkeypatch and pytest-mock, useful for
automatically cleaning up after tests and isolating the software under test from other parts of the code.
Including these instructions improves the generated tests; nonetheless
we still observe test generation that includes side effects.
@@
Within the excerpt in part 6, CoverUp tags certain lines with
their numbers, as in in “19:” in Figure 3.
It tags all lines lacking coverage as well as all lines
that are part of a branch that lacks coverage.
This tagging improves the LLM’s understanding of the missing coverage:
we find that prompts that only indicate the starting line number at the beginning of the
excerpt lead to more tests being generated that do not improve coverage.
@@ 3.3. Verification and Continued Chat
Once the LLM generates tests in response to the initial prompt, CoverUp executes them,
again measuring coverage.
In our implementation, this step is made more efficient by SlipCover’s near-zero overhead:
using coverage.py, the only other alternative tool for Python,
introduces up 260% overhead (Altmayer Pizzorno and Berger, 2023).
@@
If they do not increase coverage, or if they result in failures or errors,
CoverUp continues the chat session, pointing out the problem(s) and requesting
improvements.
Figures 4 and 5 show examples
of these prompts.
@@
requests a repair (2) and includes an excerpt of the error messages (3).
In the original execution from which this example is taken,
the LLM responds with an usable test (not shown).
Prior to executing a new test, CoverUp looks for any Python modules used by
the test that are absent from the system.
These missing modules are typically test helper modules, such as the pytest-ansible plugin
used to help test the ansible package;
Our implementation offers options to automatically install missing modules and to
record these in Python’s requirements.txt format to facilitate making
their use in a subsequent run.
@@ 3.4. Tool Functions
CoverUp exposes a get_info tool function that the LLM
can call to request additional information about any names in the excerpt,
such as functions, classes, variables or modules.
It responds to such function calls by performing static analysis and slicing relevant
portions of the source code.
Figure 6. Example of a prompt:
In response to this initial prompt, shown abbreviated, the LLM might
respond with a get_info("code.A") function call, as requires
more information on how to instantiate A in tests.
@@
Asked about code.A, CoverUp would perform static analysis and,
discovering that A is a class, respond with a slice of its definition.
In this case, it includes its __init__ constructor, which indicates to
the LLM that A’s constructor takes the initial_x keyword
argument.
The slice does not include any of the contents of method m; the LLM could
subsequently ask for code.A.m to obtain more information on it.
For example, when given the prompt shown in Figure 6, the LLM
might ask for information on code.A.
The function response, shown in Figure 7, indicates that values
of x must be passed to the class constructor using a keyword argument.
This response lets the LLM immediately respond with correct tests; we find that if we do not
include the tool function in this situaiton, the LLM attempts to simply pass values of x to the
constructor, leading to errors.
CoverUp makes get_info available both with initial (Section 3.2),
and continuation prompts (Section 3.3), so that it can help generate
both better responses and better corrections.
@@ 3.5. Integration Check
While CoverUp only saves tests that pass when ran by themselves, it is still
possible that some of these do not clean up properly, or otherwise have
side effects (“state pollution”) that cause other tests to fail (Gyori et al., 2015).
This issue is not unique to LLM-based test generation: Fraser et al. (Fraser and Arcuri, 2014)
report side effects on 50% of test classes generated by EvoSuite (Fraser and Arcuri, 2011),
an SBST test generator.
Figure 8 shows an example where a generated test
not only modifies, but also, in a misguided attempt to clean up, deletes global
symbols used by the code under test.
For a test generation approach to be effective, it must thus include some way to
handle state pollution.
@@
This approach prevents memory-based state
pollution by one test from affecting other tests.
CoverUp implements efficient test isolation based on the Unix fork()
system call for the pytest framework.
This option is the default, which Section 4 uses
for the empirical evaluation of CoverUp.
@@
This approach executes the entire suite and, upon a test failure,
searches for the polluting tests and disables them.
CoverUp implements the search by successively reducing and executing subsets
of the entire suite;
@@
This approach may lead to a significant loss in coverage as tests may
be disabled unnecessarily, and may still leave a polluting test enabled,
but allows the user to move on quickly to prompting for more tests.
@@
This LLM-generated test overwrites global constants and then,
in a misguided attempt to clean up, deletes them.
The test succeeds if executed by itself, but when executed along with
other tests, the missing constants cause other tests to fail.
Our implementation offers three alternatives to handle these
situation: it can execute tests in isolation (the default), so that
and disable the polluting tests, or disable the failing tests.
@@ 3.6. Handling Flaky Tests
Flaky tests are those that can both pass or fail inconsistently,
without any changes to the code under test (Parry et al., 2021; Lam et al., 2020; Gruber et al., 2024).
Their unreliable behavior has a variety of causes;
in the context of LLM generated tests, they can also in part
result from not providing the LLM sufficient information.
Figure 9 shows a test for the
mimesis data generator package.
The code under test randomly assigns a blood type to its
Person object; the LLM, not having been provided with
the list of valid names, assumes that it includes 0-
and 0+, when the actually code uses O-
and O+ (using the letter O, not zero).
@@
This test checks that the Person’s
blood type, randomly assigned, is valid.
It fails whenever the blood type is O+ or O-,
as the LLM assumed these would be named using zero, rather than
the letter O.
CoverUp executes each test a few times, making it more likely to find
the error and give the LLM a chance to correct it.
Our implementation uses the pytest-repeat module to execute each
newly generated test multiple times, making any flaky tests more likely to fail.
If a test fails, CoverUp continues the chat, giving the LLM a chance
to correct the problem.
@@ 3.7. Other Technical Challenges
Each individual prompt sent through OpenAI’s API typically requires several seconds
to complete.
Similarly, executing individual LLM-generated tests and measuring the new coverage
achieved requires additional time.
CoverUp repeats this process for each code segment as well as
each time the dialogue is continued.
Even though our implementation limits the amount of time spent waiting on responses and
test executions, if each code segment was processed serially, it would
take an unacceptable amount of time to create tests for packages of nontrivial size.
Instead, it prompts for tests and verifies them asynchronously, using the Python
asyncio package.
As a result, during CoverUp’s evaluation (Section 4) we observe
a 500x speedup over serial execution.
Other practical challenges arise because OpenAI’s API is
provided as a cloud service and subject to various limits.
Our implementation handles a number of timeout and other error conditions automatically.
To avoid exceeding the rate limits imposed by OpenAI, it spreads out its
requests using a leaky bucket scheme (Turner, 1986) implemented by the
Python module aiolimiter.
Additionally, we implement checkpointing to files, allowing the user
to resume prompting for tests (and not lose any progress so far) after
interrupting it or stopping due to unforeseen circumstances, such as the
OpenAI account running out of funds.
@@ 4.1. Experimental Setup
Originally collated by the authors of CodaMosa (Lemieux et al., 2023) and available
from https://github.com/microsoft/codamosa, it is derived from 35 open source
projects used in the evaluation of BugsInPy (Widyasari et al., 2020) and
Pynguin (Lukasczyk et al., 2023).
@@
PY, a set of modules originally excluded from CodaMosa’s suite because
Pynguin already performs well on it.
We evaluate CoverUp on these modules so as not to leave open the question of
how well it performs on code where Pynguin / SBST already performed well.
@@
MT, a dataset of functions extracted from the HumanEval dataset (Chen et al., 2021),
originally used in the evaluation of MuTAP (Dakhel et al., 2024).
We use this dataset to enable comparisons to MuTAP, since its implementation
lacks the support needed to to run on arbitrary Python packages.
@@
CM excludes the mimesis, sanic and thef*ck packages
from the original CodaMosa suite: mimesis is a package for generating random
data, making it difficult to generate non-flaky tests for it.
mimesis was included in CodaMosa’s original evaluation, but the tests CodaMosa
generates do not contain any assertions, and thus do not fail in the face of randomly generated values.
The sanic and thef*ck packages require modules that either are no longer
available or conflict with CoverUp.
@@
To examine RQ1, we compare against two versions of CodaMosa and
four versions of MuTAP, varying their LLM and prompt type:
@@
Since Codex is no longer available, we use original tests from CodaMosa’s
evaluation, using its best performing configuration (0.8-uninterp),
available from https://github.com/microsoft/codamosa-dataset;
CodaMosa (gpt4o), our adapted version that uses the same model as
CoverUp, to help rule out differences in performance due to the use of
different LLMs.
MuTAP (codex few-shot), the original version that uses the Codex LLM,
using a “few-shot” prompt, where two examples of test generation are included
with the prompt.
Since Codex is no longer available, we use original tests from MuTAP’s
evaluation, available from https://github.com/ExpertiseModel/MuTAP;
@@
MuTAP (gpt4o few-shot), our adapted version that uses the same model
as CoverUp, with newly generated tests using a “few-shot” prompt;
MuTAP (gpt4o zero-shot), our adapted version, but using a “zero-shot”
prompt.
To create the gpt4o versions of CodaMosa and MuTAP, we modify these to use
OpenAI’s chat API and, in the case of CodaMosa, insert instructions requesting
a code completion before its original code completion prompt.
For MuTAP, we insert a prompt requesting that it respond in the same format as
a Codex model.
We were unable to empirically evaluate (Ryan et al., 2024) as at the time of writing
it remains unavailable: its authors have indicated that it is not yet publicly available.
For RQ2, we create an “CoverUp (ablated)”, a version of CoverUp which utilizes
a nearly identical prompt, except in that it does not specify how the code lacks
coverage or tags lines lacking coverage, does not include computed import
statements, does not offer a tool function for additional context, and does not
continue the chat in case of errors or lack of coverage.
@@
We utilize the line, branch and combined line and branch coverage as metrics,
computing these over all benchmark modules.
@@
It is necessary to include both line and branch coverage as metrics because
in Python, branch coverage does not subsume line coverage:
various situations can lead the Python interpreter to throw exceptions,
and when thrown, these exceptions are not recorded as branches in coverage
information.
We run all generated tests in isolation (see Section 3.5)
and measure all coverage using SlipCover (Altmayer Pizzorno and Berger, 2023).
@@
We evaluate both CodaMosa and CoverUp using the codamosa-docker
Docker image available at https://github.com/microsoft/codamosa, modified only
to install SlipCover and to disable its default entrypoint script, allowing easy
execution of other scripts.
@@
Prior to each measurement, our scripts install the benchmark-specific package.txt
requirements file distributed with CodaMosa using pip.
Unfortunately, pip fails to install a number of these requirements.
Since these failures were originally ignored for CodaMosa, we ignore them as well,
in an effort to replicate the original conditions as closely as possible.
For MuTAP and the tests it generates, we use Python 3.9.12: MuTAP uses MutPy (Hałas, 2019)
for mutation testing, which requires older versions of Python.
@@
We configure CoverUp to use OpenAI’s gpt-4o-2024-05-13 LLM,
setting its “temperature” to zero, and do not limit the number of output tokens.
We leave CoverUp’s target code segment size (see Section 3.1)
at its default of 50 lines, and automatically repeat test executions
to look for flaky tests (see Section 3.6).
To place CoverUp on the same footing as CodaMosa, we first run it without any
pre-existing test suite.
We then run it two more times, allowing it to build upon the test suite from the
previous runs.
Since GPT-generated tests often assume the presence of certain Python modules,
we configure CoverUp to automatically install any such modules if missing.
4.2. Comparison to the previous state of the art
To compare CoverUp to the previous state of the art, we first evaluate it on the CM
suite, using CodaMosa (codex) and CodaMosa (gpt4o) as baselines.
Figure 10 shows, on the left, the line, branch and combined line and
branch coverage obtained by CoverUp and the baselines over the entire benchmark suite;
on the right, it shows the median of those measurements on a module-by-module basis.
Additionally, Figure 12 plots the difference between the
combined line and branch coverage obtained by CoverUp and CodaMosa on that suite,
also on a module-by-module basis;
green bars in the plot indicate modules where CoverUp achieved higher coverage.
As the figures show, CoverUp achieves higher line, branch and combined
line and branch coverage than both CodaMosa baselines, both measuring over the entire
benchmark code base and on a per-module basis.
Across the entire benchmark suite, CoverUp achieves 64% (vs. 54% and 51%) line coverage,
49% (vs. 34% and 27%) branch coverage, and 60% (vs. 49% and 45%) line+branch coverage;
on a per-module basis, CoverUp achieves 82% (vs. 63% and 54%) line coverage,
67% (vs. 36% and 19%) branch coverage, and 80% (vs. 56% and 47%) line+branch coverage.
We also observe that CodaMosa (gpt4o)’s performance falls slightly behind that of
CodaMosa (codex), with overall coverage measurements within 5% of each other.
@@ 4.2. Comparison to the previous state of the art
We next compare CoverUp to MuTAP, evaluating it on the MT suite.
Figure 11 shows, on the left, the line, branch and combined line and
branch coverage obtained by CoverUp and the baselines over the entire benchmark suite;
on the right, it shows the median of those measurements on a module-by-module basis.
As the figures show, CoverUp achieves higher line, branch and combined line and branch
coverage than the MuTAP baselines when measuring over the entire benchmark,
and greater or equal coverage when measuring on a per-module basis.
Across the entire benchmark suite, CoverUp achieves 90% (vs. 82% to 16%) line coverage,
89% (vs. 73% to 0%) branch coverage, and 90% (vs. 79% to 10%) line+branch coverage;
on a per-module basis, like its baselines, CoverUp achieves 100% on all metrics,
except for MuTAP (gtp4o few-shot), which achieves extremely low coverage
(17%, 0% and 11%).
We investigate MuTAP (gpt4o few-shot)’s extremely low performance and discover that
its original few-shot prompt confuses GPT-4o, leading it to generate tests for
the examples that make up the “few shot”, rather than for the function under test.
When this happens, even if the resulting tests run to completion, they do not
contribute to coverage.
As the consistently high per-module median coverage values indicate, the MT suite does not
contain particularly challenging code.
In fact, the functions in the suite are entirely self contained, only rarely utilize external code,
and when they do so, from standard libraries, and commonly include type annotations.

### v3

@@ 3.2. Initial Prompting
Figure 3 shows an example of an initial prompt, with circled numbers identifying sections.
@@
a statement assigning the LLM the persona (White et al., 2023) of an “expert Python test-driven developer”, intended to help guide it towards high quality tests;
a sentence pointing out the code excerpt (segment), identifying what file it comes from, and stating what lines or branches do not execute when tested.
The portion specifying the lines and branches missing coverage is compressed using line ranges, simplifying the prompt and reducing token usage.
a request for pytest test functions and an encouragement for the LLM to use the provided tool function;
a series of other requests, such as “include assertions” and “avoid state pollution”, to steer the result towards usable tests;
a request that the response only include the new Python tests to facilitate its extraction from the response and to reduce token usage; and
@@
We find that tests generated by GPT-4 often include top-level code calling into pytest.main or into parts of the test itself.
While such top-level code can make sense in a standalone test file, Python executes it as part of the loading process, and doing so may significantly disrupt pytest’s operation.
In fact, in some of these early test generations, such calls caused pytest to restart its test discovery, slowing it down until it became unusable.
For that reason, part 4 of the prompt directs the LLM not to include such calls.
Prompt part 4 also instructs the LLM to generate tests that “clean up […], so as to avoid state pollution”.
It suggests using monkeypatch and pytest-mock, useful for automatically cleaning up after tests and isolating the software under test from other parts of the code.
Including these instructions improves the generated tests; nonetheless, we still observe test generation that includes side effects.
@@
Prior to the code excerpt in part 6, CoverUp adds import statements that provide context for each code segment: we find that without these, LLMs often make incorrect assumptions about symbols in the code, leading to errors.
Rather than extract these verbatim from source code, CoverUp generates imports based on static analysis.
In particular, it converts any relative imports to absolute imports.
Using relative imports in code excerpts can lead to tests that use them outside of package scope, where they are invalid.
Within the code excerpt, CoverUp tags certain lines with their numbers, as in “19:” in Figure 3.
It tags all lines lacking coverage and those that are part of a branch lacking coverage.
This tagging improves the LLM’s understanding of the missing coverage: we find that prompts that only indicate the starting line number at the beginning of the excerpt lead to tests that do not improve coverage.
To send the prompt using OpenAI’s API, CoverUp embeds it as a “message” in a JSON-formatted request that also includes other fields specifying the model to use, meta-parameters such as the model temperature, etc.
Figure 4 shows an example.
Figure 4. Sending a prompt: to prompt and LLM using OpenAI’s chat API, CoverUp puts together a JSON-formatted request that includes the prompt in “messages”.
[...] indicates a portion omitted for brevity.
@@ 3.3. Verification and Continued Chat
Once the LLM generates tests in response to the initial prompt, CoverUp executes them, again measuring coverage.
In our implementation, this step is made more efficient by SlipCover’s near-zero overhead: using coverage.py, the only other alternative tool for Python, introduces up to 260% overhead (Altmayer Pizzorno and Berger, 2023).
@@
If, conversely, they do not increase coverage or result in failures or errors, CoverUp continues the chat session, pointing out the problem(s) and requesting improvements.
To continue the chat, CoverUp sends another request to the LLM, appending the LLM’s response and a new prompt to the previous messages.
Figure 5 shows how CoverUp continues the chat; Figures 6 and 7 show examples of prompts requesting improvements.
Figure 5. Continuing the chat: to continue a chat, CoverUp sends out a new request, including the previous messages (here, the initial prompt), the LLM’s response (indicated with the “assistant” role), and the new prompt.
[...] indicates a portion omitted for brevity.
@@
In the original execution from which this example is taken, the LLM responds with a usable test (not shown).
Before executing a new set of tests, CoverUp looks for any Python modules used by the test that are absent from the system.
These missing modules are typically test helper modules, such as the pytest-ansible plugin used to help test the ansible package.
Our implementation offers options to install missing modules automatically and record these in Python’s requirements.txt format to facilitate their use in a subsequent run.
@@ 3.4. Tool Functions
CoverUp exposes a get_info tool function, which allows the LLM to request additional information about any names in the excerpt, such as types or variables.
In response, CoverUp provides a portion of the source code that shows the definition of the requested object.
To indicate support for the get_info tool function, CoverUp includes its description in a “tools” JSON element within each chat request.
If the LLM needs to invoke this function, it may respond with a list of get_info call requests instead of generating text, each specifying the symbol to retrieve.
CoverUp then continues the chat by sending a new request that appends the LLM’s call requests and the retrieved results from get_info to the previous messages.
Figure 8. Example of a prompt likely to lead to a tool function call:
In response to this initial prompt, the LLM would likely choose to respond with a request to call get_info("code.A"), as it requires more information on how to instantiate A in tests.
[...] indicates a portion omitted for brevity.
@@
Asked about code.A, CoverUp performs static analysis and, discovering that A is a class, responds with an abbreviated form of its definition.
In this case, it includes its __init__ constructor, which indicates to the LLM that A’s constructor takes the initial_x keyword argument.
The response does not include any of the contents of method x; the LLM could subsequently ask for code.A.x to obtain more information on it.
The ... in the Python excerpt is literal: it is a valid no-op token in Python, and CoverUp uses it to indicate where code has been removed.
For example, when given the prompt shown in Figure 8, the LLM might ask for information on code.A.
The function response, shown in Figure 9, indicates that values of x must be passed to the class constructor using a keyword argument.
This response lets the LLM immediately respond with correct tests; we find that if we do not include the tool function in this situation, the LLM attempts to pass values of x to the constructor, leading to errors.
To generate its response, get_info first looks for a function, class, module, or variable matching the requested name.
Supporting variables is particularly important, as constants and aliases in Python are created using variable assignments.
Beginning with the module where the excerpt originated, get_info may extend its search to other modules by following import statements.
Once it locates the definition, get_info constructs a response by extracting relevant portions from each module.
For many constructs, such as assignments, import statements, or function definitions, it includes the entire definition.
However, for classes and modules, including their full definitions may result in prohibitively long responses.
Simply omitting elements, as done in the initial prompt, can lead to missing essential information or cause the LLM to assume such elements do not exist.
For example, if get_info had entirely omitted property x in Figure 9, there would be no indication of how attribute A.x was accessed.
Instead, get_info replaces the bodies of such elements with ..., a valid “no-op” token in Python, and indicates the omission in the response.
CoverUp makes get_info available both with initial (Section 3.2) and continuation prompts (Section 3.3) so that it can help generate both better responses and better corrections.
@@ 3.5. Integration Check
While CoverUp only saves tests that pass when run in isolation, it is still possible that some may fail to clean up after themselves or introduce side effects (“state pollution”) that cause other tests to fail (Gyori et al., 2015).
This issue is not unique to LLM-based test generation: Fraser et al. (Fraser and Arcuri, 2014) report side effects on 50% of test classes generated by EvoSuite (Fraser and Arcuri, 2011), an SBST test generator.
Figure 10 shows an example where a generated test not only modifies but also deletes global symbols used by the code under test, in a misguided attempt to clean up.
For a test generation approach to be effective, it must thus include some way to handle state pollution.
@@
This approach prevents memory-based state pollution by one test from affecting other tests.
CoverUp implements efficient test isolation based on the Unix fork() system call for the pytest framework.
This option is the default, which Section 4 uses for the empirical evaluation of CoverUp.
@@
This approach executes the entire suite and, upon a test failure, searches for the polluting tests and disables them.
CoverUp implements the search by successively reducing and executing subsets of the entire suite;
@@
This approach may lead to a significant loss in coverage as tests may be disabled unnecessarily and may still leave a polluting test enabled, but it allows the user to move on quickly to prompting for more tests.
@@
This LLM-generated test overwrites global constants and then deletes them in a misguided attempt to clean up.
The test succeeds if executed by itself, but when executed along with other tests, the missing constants cause other tests to fail.
Our implementation offers three alternatives to handle this situation: it can execute tests in isolation (the default), disable the polluting tests, or disable the failing tests.
[...] indicates a portion omitted for brevity.
@@ 3.6. Handling Flaky Tests
Flaky tests are those that can both pass or fail inconsistently without any changes to the code under test (Parry et al., 2021; Lam et al., 2020; Gruber et al., 2024).
Their unreliable behavior has a variety of causes; in the context of LLM-generated tests, they can also, in part, result from not providing the LLM with sufficient information.
Figure 11 shows a test for the mimesis data generator package.
The code under test randomly assigns a blood type to its Person object; the LLM, not having been provided with the list of valid names, assumes that it includes 0- and 0+ whereas code actually uses O- and O+ (using the letter O, not zero).
@@
This test checks that a randomly assigned blood type is valid.
It fails whenever the blood type is O+ or O-, as the LLM assumed these would be named using zero rather than the letter O.
CoverUp executes each test a few times, making the error more likely to surface and thus allowing the LLM to attempt to correct it.
Our implementation uses the pytest-repeat module to execute each newly generated test multiple times, making any flaky tests more likely to fail.
If a test fails, CoverUp continues the chat, allowing the LLM to attempt to correct the problem.
@@ 3.7. Other Technical Challenges
Each individual prompt sent through OpenAI’s API typically requires several seconds to complete.
Similarly, executing individual LLM-generated tests and measuring the new coverage achieved requires additional time.
CoverUp repeats this process for each code segment and each time the dialogue is continued.
Even though our implementation limits the time spent waiting on responses and test executions, if each code segment were processed serially, creating tests for packages of nontrivial size would take an unacceptable amount of time.
Instead, it prompts for tests and verifies them asynchronously, using the Python asyncio package.
As a result, during CoverUp’s evaluation (Section 4), we observe a 500x speedup over serial execution.
Other practical challenges arise because OpenAI’s API is provided as a cloud service and is subject to various limits.
Our implementation handles various timeout and other error conditions automatically.
To avoid exceeding the rate limits imposed by OpenAI, it spreads out its requests using a leaky bucket scheme (Turner, 1986) implemented by the Python module aiolimiter.
Additionally, we implement checkpointing to files, allowing the user to resume prompting for tests (and not lose any progress so far) after interrupting it or stopping due to unforeseen circumstances, such as the OpenAI account running out of funds.
@@ 4. Evaluation
RQ4::
How does the cost of running CoverUp compare to CodaMosa?
(Section 4.5)
RQ5::
How important are CoverUp’s components to its performance?
(Section 4.6)
@@ 4.1. Experimental Setup
Collated originally by the authors of CodaMosa (Lemieux et al., 2023) and available from https://github.com/microsoft/codamosa, it is derived from 35 open-source projects used in the evaluation of BugsInPy (Widyasari et al., 2020) and Pynguin (Lukasczyk et al., 2023).
@@
PY, a set of modules originally excluded from CodaMosa’s suite because Pynguin already performs well on it.
We evaluate CoverUp on these modules so as not to leave open the question of how well it performs on code where Pynguin / SBST already performed well.
@@
MT, a dataset of functions extracted from the HumanEval dataset (Chen et al., 2021), originally used in the evaluation of MuTAP (Dakhel et al., 2024).
We use this dataset to enable comparisons to MuTAP since its implementation lacks the support needed to run on arbitrary Python packages.
@@
CM excludes the mimesis, sanic, and thef*ck packages from the original CodaMosa suite: mimesis is a package for generating random data, making it challenging to generate non-flaky tests for it.
mimesis was included in CodaMosa’s original evaluation, but the tests CodaMosa generates do not contain any assertions and thus do not fail in the face of randomly generated values.
The sanic and thef*ck packages require modules that either are no longer available or conflict with CoverUp.
@@
To examine RQ1, we compare against two versions of CodaMosa and four versions of MuTAP, varying their LLM and prompt type:
@@
Since Codex is no longer available, we use original tests from CodaMosa’s evaluation, using its best performing configuration (0.8-uninterp), available from https://github.com/microsoft/codamosa-dataset;
CodaMosa (gpt4o), our adapted version that uses the same model as CoverUp, to help rule out differences in performance due to the use of different LLMs.
MuTAP (codex few-shot), the original version that uses the Codex LLM, using a “few-shot” prompt, where two examples of test generation are included with the prompt.
Since Codex is no longer available, we use original tests from MuTAP’s evaluation, available from https://github.com/ExpertiseModel/MuTAP;
@@
MuTAP (gpt4o few-shot), our adapted version that uses the same model as CoverUp, with newly generated tests using a “few-shot” prompt;
MuTAP (gpt4o zero-shot), our adapted version, but using a “zero-shot” prompt.
To create the gpt4o versions of CodaMosa and MuTAP, we modify these to use OpenAI’s chat API and, in the case of CodaMosa, insert instructions requesting a code completion before its original code completion prompt.
For MuTAP, we insert a prompt requesting that it respond in the same format as a Codex model.
We could not empirically evaluate (Ryan et al., 2024) as, at the time of writing, it remains unavailable: its authors have indicated that it is not publicly available.
For RQ2, we create an “CoverUp (ablated)”, a version of CoverUp which utilizes a nearly identical prompt, except in that it does not specify how the code lacks coverage or tags lines lacking coverage, does not add computed import statements, does not offer a tool function for additional context, and does not continue the chat in case of errors or lack of coverage.
@@
We utilize the line, branch, and combined line and branch coverage as metrics, computing these for all benchmark modules.
@@
It is necessary to include both line and branch coverage as metrics because in Python, branch coverage does not subsume line coverage: various situations can lead the Python interpreter to throw exceptions, and when thrown, these exceptions are not recorded as branches in coverage information.
We run all generated tests in isolation (see Section 3.5) and measure all coverage using SlipCover (Altmayer Pizzorno and Berger, 2023).
@@
We evaluate both CodaMosa and CoverUp using the codamosa-docker Docker image available at https://github.com/microsoft/codamosa, modified only to install SlipCover and to disable its default entrypoint script, allowing easy execution of other scripts.
@@
As the host system for Docker, we utilize a Linux kernel 5.6 system with 10 Intel i9 cores at 3.7GHz and 64GB RAM.
Before each measurement, our scripts install the benchmark-specific package.txt requirements file distributed with CodaMosa using pip.
Unfortunately, pip fails to install some of these requirements.
Since these failures were originally ignored for CodaMosa, we also ignore them to replicate the original conditions as closely as possible.
For MuTAP and the tests it generates, we use Python 3.9.12: MuTAP uses MutPy (Hałas, 2019) for mutation testing, which requires older versions of Python.
@@
We configure CoverUp to use OpenAI’s gpt-4o-2024-05-13 LLM, setting its “temperature” to zero and do not limit the number of output tokens.
We leave CoverUp’s target code segment size (see Section 3.1) at its default of 50 lines and automatically repeat test executions to look for flaky tests (see Section 3.6).
We first run it without any pre-existing test suite to place CoverUp on the same footing as CodaMosa.
We then run it twice more, allowing it to build upon the test suite from the previous runs.
Since GPT-generated tests often assume the presence of certain Python modules, we configure CoverUp to install any such missing modules automatically.
4.2. [RQ1] Comparison to the previous state of the art
To compare CoverUp to the previous state of the art, we first evaluate it on the CM suite, using CodaMosa (codex) and CodaMosa (gpt4o) as baselines.
Figure 12 shows, on the left, the line, branch, and combined line and branch coverage obtained by CoverUp and the baselines over the entire benchmark suite; on the right, it shows the median of those measurements on a module-by-module basis.
Additionally, Figure 14 plots the difference between the combined line and branch coverage obtained by CoverUp and CodaMosa on that suite, also on a module-by-module basis; green bars in the plot indicate modules where CoverUp achieved higher coverage.
Across the board, CoverUp yields higher coverage than CodaMosa, whether measured over the entire suite or on a module-by-module basis.
The graph shows the difference in (lines + branches) coverage between CoverUp and CodaMosa for the modules in the CM suite.
Green highlights modules where CoverUp achieved a higher coverage.
Figure 14. [RQ1] CoverUp also performs well on the PY suite:CoverUp yields 100% median and near 100% overall coverage on all metrics, showing that its effectiveness is not limited to the modules selected for the CM suite.
Figure 15. [RQ1] CoverUp yields higher overall and per-module coverage:CoverUp yields higher or equal coverage than MuTAP, whether measured over the entire suite or on a module-by-module basis.
As the figures show, CoverUp achieves higher line, branch, and combined line and branch coverage than both CodaMosa baselines, both measuring over the entire benchmark code base and on a per-module basis.
Across the entire benchmark suite, CoverUp achieves 64% (vs. 54% and 51%) line coverage, 49% (vs. 34% and 27%) branch coverage, and 60% (vs. 49% and 45%) line+branch coverage.
On a per-module basis, CoverUp achieves 82% (vs. 63% and 54%) line coverage, 67% (vs. 36% and 19%) branch coverage, and 80% (vs. 56% and 47%) line+branch coverage.
These per-module improvements over CodaMosa (GPT-4o) and CodaMosa (Codex) are statistically significant: using paired permutation tests, we obtain a -value of for both, well below the standard threshold.
We also observe that CodaMosa (gpt4o)’s performance falls slightly behind that of CodaMosa (codex), with overall coverage measurements within 5% of each other.
@@
As Figure 14 shows, CoverUp does not achieve higher coverage than
CodaMosa for some modules.
We examine CoverUp’s logs and observe that timeouts running tests for modules in the youtube_dl package constitute the largest single source of failure.
While the authors of CodaMosa also report problems with timeouts for that package (Lemieux et al., 2023), it seems likely that CoverUp’s test execution timeout of one minute was set too low.
In other cases, we observe that CoverUp’s static analysis cannot provide the LLM with correct context when the code under test contains conditional imports, such as those intended to accommodate package differences across Python versions.
Additionally, while LLM test generations often include multiple test functions in the same response, CoverUp rejects them all if any of them fails; it is possible that CoverUp’s performance would have been higher if it were to evaluate each test function individually.
Next, we evaluate the performance of CoverUp on code for which the original Pynguin performed well.
Figure 14 shows the results.
Given CoverUp’s 100% median per-module coverage and near 100% overall coverage results, we conclude that CoverUp also performs well on such code.
We then compare CoverUp to MuTAP, evaluating it on the MT suite.
Figure 15 shows, on the left, the line, branch, and combined line and branch coverage obtained by CoverUp and the baselines over the entire benchmark suite; on the right, it shows the median of those measurements on a module-by-module basis.
As the figures show, CoverUp achieves higher line, branch, and combined line and branch coverage than the MuTAP baselines when measuring over the entire benchmark and greater or equal coverage when measuring on a per-module basis.
Across the entire benchmark suite, CoverUp achieves 90% (vs. 82% to 16%) line coverage, 89% (vs. 73% to 0%) branch coverage, and 90% (vs. 79% to 10%) line+branch coverage; on a per-module basis, like its baselines, CoverUp achieves 100% on all metrics, except for MuTAP (gtp4o few-shot), which achieves extremely low coverage (17%, 0% and 11%).
We investigate MuTAP (gpt4o few-shot)’s extremely low performance and discover that its original few-shot prompt confuses GPT-4o: it generates tests for the test generation examples included in that prompt rather than for the function under test.
When this happens, even if the resulting tests run to completion, they do not contribute to coverage.
As the consistently high per-module median coverage values indicate, the MT suite does not contain particularly challenging code.
In fact, the functions in the suite are entirely self-contained.
They only rarely utilize external code; when they do so, they use it from standard libraries.
Additionally, the functions commonly include type annotations.